ON A GENERALIZATION OF MORISITA’S MODEL FOR ESTIMATING THE HABITAT PREFERENCE

CH. A. CHARALAMBIDES AND M. V. KOUTRAS

Department of Mathematics, Section of Statistics and O.R., University of Athens, Panepistemiopolis, Athens 157 84, Greece

(Received January 27, 1992; revised July 6, 1992)

Abstract. The Morisita’s model for estimating the habitat preference by the ant lions Genuroides japonicus is generalized by introducing, in addition to the environmental densities a and b, a repulsivity parameter θ. The probability function of the number L_{n} of individuals choosing fine sand to settle when a total of n ant lions are introduced is examined. A heuristic and the minimum chi-square methods for estimating the parameters a, b and θ are discussed.

Key words and phrases: Eulerian numbers, Gould and Hopper numbers, minimum chi-square estimation, unimodality.